OpenStudy (anonymous):

A screen saver displays the outline of a 3 cm by 2 cm rectangle and then expands the rectangle in such a way that the 2 cm side is exanpanding at the rate of 4 cm/sec and the proportions of the rectangle never change. How fast is the area of the rectangle increasing when its dimensions are 12 cm by 8 cm?

OpenStudy (mathmate):

This is a linearization problem. A = uv dA/dt = u (dv/dt) + v(du/dt) du/dt=4 cm/s. dv/dt =4*(3/2)=6 cm/s (aspect ratio does not change) When u=8, v=12, then dA/dt = (8*6 + 12*4) cm^2/s = 96 cm^2/s

OpenStudy (anonymous):

o thx my teacher vhasnt really explained those yet

OpenStudy (mathmate):

But was it done in a way you could understand? If not, we have to try another way of doing it.

OpenStudy (anonymous):

i dont get uv

OpenStudy (mathmate):

u, v represent the sizes of the screen at time t, i.e. u=u(t), v=v(t).